# Tokens, the oft-overlooked appetizer: Large language models, the distributional hypothesis, and meaning

### v7

@@ 3.5 Bias, alignment, and ethics
The last two points especially make up a “backdoor” for unwanted content to make its way into the model’s internal representations: as seen with the “rung” and “run” example (Fig. 40), any token that is an orthographic component of another word could retain information about that word. [^90] As long as that word was seen during relevant training periods; again, recall the apparent source and quality of training data. Since many of the prevalent tokenization strategies are popular in large part because they handle OOV words by including individual letters and small subword units, [^91] Technically, often at the byte level essentially any given word is guaranteed to have token components in the typical model vocabulary. Templeton et al., 2024 attest this composition of concepts with respect to the feature level (a higher level within the model than the token): “[i]mportantly, not having a feature dedicated to a particular concept does not mean that the reconstructed activations do not contain information about that concept, as the model can use multiple related features compositionally to reference a specific concept” Templeton et al. (2024).
@@
Beyond alignment, based on the the in-domain problem and the ouroboros: archaic, obscure, and obsolete words are likely not going to be tokenized. [^99] Our results are consistent with this finding from Zhemchuzhina et al., 2022 Zhemchuzhina, Filippov, and Yamshchikov (2022)! How is that going to contribute to loss of beautiful cultural artifacts? Beyond buildings and even texts, language itself is a heritage site.100100100For example, peruse the Phrontistery Chrisomalis (1996-2021)! Similar arguments have been made at the narrative and cultural level, beyond the individual word or phrase: Generative AI acts as a “crucible of mediocrity” Ruiz, Zimmerman, and team (2023), as a regression to the mean Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).101101101We speculate that the probability mass of the logits corresponding to the subset of tokens involved in inference is likely to be relatively high, and that this is related to why models sometimes behaviourally collapse into inane repetition. More broadly, that repetition is a symptom of the iterative concentration of future plausibility based on previous plausibility, which is effectively a one-way ratchet Zimmerman et al. (2024). Imagine a hypothetical LLM generating output based on every previous utterance, including what it, itself, has said. As an analogy, compare sequence A005150 Foundation Inc. (2024a) to sequence A363054 Foundation Inc. (2024b): the initial seeds are obviously different, but as the output length grows, those differences are essentially lost to time (history being deeper on the right side of each term), and the sequences feel increasingly indistinguishable, locked in to reiterating the same sorts of patterns. It tends towards being normative and generic, which is undercut by creativity, novelty, and diversity Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).102102102“Beyond achieving linguistic coherence, beyond resolving the ‘hallucinations’ and errors, it is the performance of the generic and the normative, sounding ‘right,’ that sells AI tools… As currently designed, generative AI tend towards being generic, centrist, normative, and banal. Hallucinations undermine that, but so do cultural and narrative diversity” Gillespie (2024). This is, of course, partly a consequence of the fuzzing up of grammaticality or frequency with truth Zimmerman et al. (2024)!103103103McCoy et al., 2024, provides support as to (1) our characterization of plausible text generation as intertwining frequency and truth and (2) the theoretical importance of being able to modulate the primacy of the plausible text generation objective function, and the limitations of our current attempts in that vein NTP vs “reasoning”, efficacy of attempts to modulate primacy of NTP: “Specifically, o1 – like previous LLMs – is sensitive to the probability of examples and tasks, performing better and requiring fewer ‘thinking tokens’ in high-probability settings than in low-probability ones. These results show that optimizing a language model for reasoning can mitigate but might not fully overcome the language model’s probability sensitivity” McCoy et al. (2024b).,104104104For more discussion of the DH, and regime 1 vs. regime 2 learning (pattern-matching vs. learning from abstractions; hearing a novel word in context vs. reading a dictionary definition), see  Zimmerman (2025).
@@
Current LLMs have trouble with unusual data, at all scales.105105105Language, too, has this tendency inalienably within it, via its fungibility. Information technologies homogenize: that is both the benefit and price of fungibility. Fungibility is inherently degrounding, and degrounding is inherently homogenizing; it entails a loss of information or meaning, but a gain in utility. See first author’s thesis for more discussion of the construction of meaning along the fungibility-grounding spectrum Zimmerman (2025). Unusual data, data in the tails of the distribution, is important, both for cultural reasons and practical reasons of model performance and collapse. This is especially worth considering in the light of proposals to use LLMs to generate synthetic data (the ouroborous).106106106Shumailov et al., 2024 found that “in learning tasks in which the tails of the underlying distribution matter, one needs access to real human-produced data” as opposed to relying on LLM-generated data. They found that “access to the original data distribution is crucial” in order to prevent model collapse in subsequent generations Shumailov et al. (2024). Then, an initial tokenization step that curtails a model’s access to the tails of the attested word use distribution by omitting longer and rare words could also contribute to model collapse or other performance issues. Tokenization is a deterministic (though context-dependent) process, from within the perspective of a single algorithm being applied to a single fixed corpus. But tokenization as a cultural artifact in a larger system is less transparent.107107107Research continues to emerge that provides additional support for this argument Yakura et al. (2025).
@@ 4 Conclusion
It is clear that we need to develop strategies that allow LLMs to make better use of the data to which they are exposed.108108108Zhuang et al., 2024 note “the immense volume of training data necessitated for effective LM performance, surpassing—by orders of magnitude—the linguistic input received during human language acquisition” Zhuang, Fedorenko, and Andreas (2024). People, by contrast, are very good at wringing new information from familiar stories.109109109Right now, models display a striking dissimilarity with people in what happens when they are repeatedly exposed to the same data, either directly, or indirectly through iterated training on synthetic data. For example, Chang et al., 2024 found that “LLMs trained with duplicated training data exhibit faster forgetting” Chang et al. (2024), and Shumailov et al., 2024, found “indiscriminate use of model-generated content in training causes irreversible defects in the resulting models, in which tails of the original content distribution disappear” Shumailov et al. (2024). On the other hand, a hallmark of human storytelling is enjoying—and getting new content out of—experiencing a story (created by other humans, even potentially by oneself) again and again. People have been studying religious texts, as in Biblical exegesis, for thousands of years, yet new ideas emerge. An individual may return to the Bible (continuing the same example) upwards of hundreds of times, and yet experience new insight. How can LLMs wring new insight from old data? This stark difference in behaviour is worth further consideration.
@@
One strategy would be to bring various representations the LLM uses into greater alignment with human structures and processes. There is evidence that humans can do quite a lot with very little data (e.g., a broad spectrum of poverty-of-the-stimulus-related arguments and processes Zhuang, Fedorenko, and Andreas (2024), including fast-mapping Gelman and Brandone (2010), and in experimental scenarios testing the distributional hypothesis McDonald and Ramscar (2001)).110110110Likewise, LLMs can sometimes do a lot with a little, such as in lexinvariant LMs Huang et al. (2023b).
@@
Furthermore, good model vocabularies may already be converging on being somewhat human-like. Symmetrically, the tokens themselves point to commonalities between the entropy-minimizing objective functions often used in tokenization and the forces that shape human language. The hundreds of thousands of years Fedorenko, Piantadosi, and Gibson (2024)111111111E. Fedorenko is an author on both Fedorenko et al.,, 2024 Fedorenko, Piantadosi, and Gibson (2024) and Zhuang et al.,, 2024 Zhuang, Fedorenko, and Andreas (2024), which seem to make arguments that are at least partially contradictory. We assume in good faith there must be a perspective (presumably E.F.’s) from which the arguments are entirely consistent, and take both papers at face value for the purposes of this project. in which only people spoke language, during which pressures of efficiency seem to have shaped its development at multiple levels, surely evolved structures we can now take advantage of in our models.
@@
We argue that the de facto approaches in each of those areas are worth revisiting.112112112The use of special tokens (e.g., SEP) in LLMs is out of scope for this paper but intriguing for future work, as is the impact of embedding vector dimension and other implementation choices. More broadly, now that the efficacy of LLM architecture has been established, we should revisit our prior underconstrained implementation choices, such as the relevant objective functions that shape the model and the functions that conjoin its components.113113113e.g., operations involving Query, Key, and Value vectors in the attention blocks, or applications of softmax.
@@
We contend that LLMs validate and isolate the distributional hypothesis. Specifically, they show the power of syntagmatic information for producing minimally viable language performance (when coupled with sufficient architecture). We hypothesize that LLMs bootstrap limited meaning, potentially including abstractions and world knowledge, from distributional patterns: more specifically, that while LLM’s start with syntagmatics, they create further structure leading to limited semantic meaning, the ceiling of which is what can be conveyed by purely distributional linguistic information and diegetic approximations composed thereof.114114114Harris’ purely linguistic meaning Sahlgren (2008). The lack of hard boundary between world knowledge and linguistic knowledge makes this feat possible but it also engenders the problems that come from fuzzing up grammaticality and truth, or frequency and truth Zimmerman et al. (2024); Gong et al. (2020),115115115Gong et al., 2020, found that in some situations, word embeddings were “biased towards word frequency”, with the embeddings of high- and low-frequency words lying in different areas of the embedding space, and the embedding of a rare word and a popular word being physically distant despite their semantic similarity. Within the embedding space, rare and popular words behaved differently, with popular words having semantically relevant neighbors, and rare words being overwhelmingly surrounded by other, semantically-unrelated, rare words Gong et al. (2020). (They foresee these structures as causing downstream performance issues, so they propose FRequency AGnostic word Embedding (FRAGE) as a solution Gong et al. (2020).) which are amplified by the choice of next token prediction as objective function.116116116And by other, more specific, architectural choices, such as how attention is implemented. These choices are related, since the attention design is in the service of generating plausible output tokens, so in the context of this paper, we are referring to the broad goal of getting the model to generate plausible tokens, not only to specific portions of the model that might be most directly technically implicated, like the decoder (if there is one), the output layer, or the final probability distribution over the vocabulary. We propose that experiential grounding enhances the depth of meaning through othermodal reference for both people and LLMs117117117Recall Wolfe and Caliskan, 2022 Wolfe and Caliskan (2022) and Zhuang et al., 2024 Zhuang, Fedorenko, and Andreas (2024). and therefore represents a promising avenue for their continued improvement.
@@
Importantly, a model’s gnogeography118118118Recall that gnogeography is the landscape of knowledge; a being’s mind-and-body, their architecture, in conjunction with what information they have taken in. can be expanded through diegetic approximations Zimmerman et al. (2024); Brown et al. (2020); Chang et al. (2024); Mahowald et al. (2024). Meaning that, with careful structuring of training data and model architecture (e.g., batch size, number of training steps), LLMs do seem to be able to acquire additional world knowledge Chang et al. (2024), implying that there is space between the world knowledge necessary for MVP language and what it is possible to know through the medium of language. Given that current LLMs clearly possess MVP language but are missing much human-familiar knowledge and common sense, they are somewhere within that space Zimmerman et al. (2024); Mahowald et al. (2024). The entirety of that space is the upper bound of their gnogeography.
@@
This is crucial to remember in the face of AI hype, one core promise of which is “we have a machine that can talk, therefore we have a machine that can think”. In fact, LLMs demonstrate the plausibility of a machine that can solely talk; while a machine may be able to both talk and think, the ‘therefore’ between those abilities is not justified.119119119“Talk” and “think” are shorthand. Of course, talking involves thinking, and thinking can be scaffolded by talking. See previous sections, especially Sec. 2.3, where we referenced stochastic parrots; we do think MVP language as implemented in LLMs involves more understanding than would be granted by that framework.
@@
The size and shape of the upper bound LLM gnogeography is unknown.120120120The hypothetical ceiling being everything that can be encoded in language. While this is exciting, it also means it is plausible that their gnogeography is incompatible with what would be required for many of the tasks we want LLMs to perform. This warning also applies to context-specific modifications to the model such as alignment and fine-tuning: strategies that do not attempt to reconcile the innate gnogeographic constraints of the model with the intended task have a flimsy foundation on which to build success. As long as generating plausible text is the goal for the model, we should expect plausible text rather than accurate text to be the output. That those two can often coincide, while alluring and often convenient, is fundamentally incidental.121121121We think this framework for thinking about LLMs can be helpful to people trying to decide how to use and study them. If an LLM is deployed as part of e.g., a transcription pipeline, the potential for unwanted hallucination can and should be assumed by default Koenecke et al. (2024), unless part of the pipeline’s design is specifically intended to counteract not just the symptomatic output but the causes - and even then, caution is necessary.
5 Post-Script
@@
Tokens of the kind discussed here may soon be on the way out as an architectural choice, at least for some AI applications. They hardwire a degree of cognition that might be better left to the model itself; historically, more open-ended architectures have often yielded performance gains Zimmerman (2025). As previously discussed, tokens represent a form of instantiated linguistic relativism – a commitment to the primacy of language for thought. This applies not only to their role as semantic primitives and as the units determining what distributional information is perceptible to the model, but also to how they function in Chain-of-Thought (CoT) reasoning,122122122What exactly ‘reasoning’ models are doing – whether via CoT or other means – remains an open question. We believe there are compelling reasons to think the gnogeographic ceiling of deep learning could be far higher than that attested by current prominent LLMs Zimmerman (2025). However, model behavior is notoriously difficult to interpret, so caution is warranted. For concision, here we assume CoT induces something beyond MVP linguistic performance – a substantive rather than stylistic difference. However, it is entirely possible that ‘reasoning’ LLMs may more aptly be described as reweighting statistical patterns under varying conditions rather than truly reasoning about new tasks. and to their relationship with the allocation of computational resources during inference (e.g., the correspondence between a single forward pass and a single token) Gladstone et al. (2025); Wang et al. (2025).
Regardless of their architectural suitability, tokens remain invaluable to linguists: they are configurable, language-related structures generated by the most performant computational models of language yet developed.123123123At least among well-known models; niche architectures may outperform LLMs on specific measures. For popular, widely used model families, however, the description holds. We plan to explore their use in linguistics further in future work (and hope others will do the same).
@@
The goal we give to the model in NTP is to create plausible text, based on text that’s been seen before (i.e. frequency as truth). As aforementioned, we are baffled as to why we would expect that to lead only to accurate text. The better-framed question (than ‘why do models hallucinate’) in our opinion is why does something that looks like reasoning ever fall out of this set up in the first place? Can we really sometimes extend the local task of predicting the next token into something like thinking?
@@
We speculate that as the scope of knowledge required by LLM tasks increases, so too will the scope of linguistic structures the models must rely upon. NTP more-or-less suffices for human-like text generation. However, with the shift toward ‘reasoning’ – e.g. the adoption of CoT prompting – the scope of the task expands beyond MVP performance (grammaticality, acceptability, morphosyntactic and inferential lexical competence) towards solving multipart logical problems invoking narrative, social rules, extensive world knowledge, and specialized techniques. Consequently, effective CoT prompting engages larger linguistic structures than NTP alone.124124124We view CoT as a (possible) extension of the extremely local next-token-prediction (NTP) objective – broadening the model-perceptible range of the knowledge expressible through language (re ascertaining the ceiling of LLM gnogeography). The NTP task more-or-less suffices for acquiring morphology, syntax, and some degree of pragmatics – sufficient to support a wide range of human-like responses. Caveat: these are not perfectly human-like competencies, as evidenced by model collapse when models are trained on their own output Shumailov et al. (2024). For pure LLMs, as we have argued herein, language has been inherently intertwined with thought via tokens – a paradigm leveraged in CoT to expand model capacity. As tasks demand more from the information encoded in language, models must become increasingly aware of linguistic structures not salient within NTP alone (at least generally, many of these structures would seem to be larger than that local scope). Thus, as ‘reasoning’ LLMs improve, they increasingly demonstrate the feasibility of language as a cognitive foundation, one strategy for how it is possible to think. Caveats remain: the boundary between world knowledge and linguistic knowledge is unclear, perhaps non-existent, and NTP may foster some non-local abstractions Dumas et al. (2025); Cui et al. (2024); Zimmerman (2025). Relationships among task design, implementation, model capacity, and data remain only loosely understood – but in broad strokes, this is the emerging picture.
@@
In our interpretation, as computer science researchers refine what they want LLMs to learn from and do with language, they are rediscovering patterns from pragmatics. For instance, the CoT principles of efficacy, efficiency, brevity, and sufficiency echo Grice’s maxims of quantity, quality, relation, and manner Cheng et al. (2025); Grice (1989). Similarly, researchers found they could incentivize greater model ‘truthfulness’125125125A loaded concept mostly out of scope for this discussion; we note ‘truth’ is – at the very least – unlikely to be a property found within the four corners of a text, loosely paraphrasing Joseph Weizenbaum. through teaching the model to abstain when it was uncertain Wei et al. (2025), clearly a Gricean approach. Re the interplay between objective function, scope, and salient linguistic structures, Grice’s maxims emerge only at higher linguistic scales (multi-utterance or conversational levels) rather than at the local level of phonemes and morphemes, and perhaps are only beginning to show up even at the level of individual utterances. We plan to continue investigating this interplay – between the DH as a fundamental learning strategy, model objectives, intended tasks, locality/ scale, and model cognition – in future work. These resonances also demonstrate that computer scientists could look to linguistics (and psychology and cognitive science) for inspiration, and in particular, for strategies which could enhance model performance (and vice versa).

### v8

@@ 3.5 Bias, alignment, and ethics
The last two points especially make up a “backdoor” for unwanted content to make its way into the model’s internal representations: as seen with the “rung” and “run” example (Figure 40), any token that is an orthographic component of another word could retain information about that word. [^91] As long as that word was seen during relevant training periods; again, recall the apparent source and quality of training data. Since many of the prevalent tokenization strategies are popular in large part because they handle OOV words by including individual letters and small subword units, [^92] Technically, often at the byte level essentially any given word is guaranteed to have token components in the typical model vocabulary. Templeton et al., 2024 attest this composition of concepts with respect to the feature level (a higher level within the model than the token): “[i]mportantly, not having a feature dedicated to a particular concept does not mean that the reconstructed activations do not contain information about that concept, as the model can use multiple related features compositionally to reference a specific concept” Templeton et al. (2024).
@@
Beyond alignment, based on the the in-domain problem and the ouroboros: archaic, obscure, and obsolete words are likely not going to be tokenized.100100100Our results are consistent with this finding from Zhemchuzhina et al., 2022 Zhemchuzhina, Filippov, and Yamshchikov (2022)! How is that going to contribute to loss of beautiful cultural artifacts? Beyond buildings and even texts, language itself is a heritage site.101101101For example, peruse the Phrontistery Chrisomalis (1996-2021)! Similar arguments have been made at the narrative and cultural level, beyond the individual word or phrase: Generative AI acts as a “crucible of mediocrity” Ruiz, Zimmerman, and team (2023), as a regression to the mean Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).102102102We speculate that the probability mass of the logits corresponding to the subset of tokens involved in inference is likely to be relatively high, and that this is related to why models sometimes behaviourally collapse into inane repetition. More broadly, that repetition is a symptom of the iterative concentration of future plausibility based on previous plausibility, which is effectively a one-way ratchet Zimmerman et al. (2024). Imagine a hypothetical LLM generating output based on every previous utterance, including what it, itself, has said. As an analogy, compare sequence A005150 Foundation Inc. (2024a) to sequence A363054 Foundation Inc. (2024b): the initial seeds are obviously different, but as the output length grows, those differences are essentially lost to time (history being deeper on the right side of each term), and the sequences feel increasingly indistinguishable, locked in to reiterating the same sorts of patterns. It tends towards being normative and generic, which is undercut by creativity, novelty, and diversity Gillespie (2024); Ruiz, Zimmerman, and team (2023); Lovato et al. (2024).103103103“Beyond achieving linguistic coherence, beyond resolving the ‘hallucinations’ and errors, it is the performance of the generic and the normative, sounding ‘right,’ that sells AI tools… As currently designed, generative AI tend towards being generic, centrist, normative, and banal. Hallucinations undermine that, but so do cultural and narrative diversity” Gillespie (2024). This is, of course, partly a consequence of the fuzzing up of grammaticality or frequency with truth Zimmerman et al. (2024)!104104104McCoy et al., 2024, provides support as to (1) our characterization of plausible text generation as intertwining frequency and truth and (2) the theoretical importance of being able to modulate the primacy of the plausible text generation objective function, and the limitations of our current attempts in that vein NTP vs “reasoning”, efficacy of attempts to modulate primacy of NTP: “Specifically, o1 – like previous LLMs – is sensitive to the probability of examples and tasks, performing better and requiring fewer ‘thinking tokens’ in high-probability settings than in low-probability ones. These results show that optimizing a language model for reasoning can mitigate but might not fully overcome the language model’s probability sensitivity” McCoy et al. (2024b).,105105105For more discussion of the DH, and regime 1 vs. regime 2 learning (pattern-matching vs. learning from abstractions; hearing a novel word in context vs. reading a dictionary definition), see  Zimmerman (2025).
@@
Current LLMs have trouble with unusual data, at all scales.106106106Language, too, has this tendency inalienably within it, via its fungibility. Information technologies homogenize: that is both the benefit and price of fungibility. Fungibility is inherently degrounding, and degrounding is inherently homogenizing; it entails a loss of information or meaning, but a gain in utility. See first author’s thesis for more discussion of the construction of meaning along the fungibility-grounding spectrum Zimmerman (2025). Unusual data, data in the tails of the distribution, is important, both for cultural reasons and practical reasons of model performance and collapse. This is especially worth considering in the light of proposals to use LLMs to generate synthetic data (the ouroborous).107107107Shumailov et al., 2024 found that “in learning tasks in which the tails of the underlying distribution matter, one needs access to real human-produced data” as opposed to relying on LLM-generated data. They found that “access to the original data distribution is crucial” in order to prevent model collapse in subsequent generations Shumailov et al. (2024). Then, an initial tokenization step that curtails a model’s access to the tails of the attested word use distribution by omitting longer and rare words could also contribute to model collapse or other performance issues. Tokenization is a deterministic (though context-dependent) process, from within the perspective of a single algorithm being applied to a single fixed corpus. But tokenization as a cultural artifact in a larger system is less transparent.108108108Research continues to emerge that provides additional support for this argument Yakura et al. (2025).
@@ 4 Conclusion
It is clear that we need to develop strategies that allow LLMs to make better use of the data to which they are exposed.109109109Zhuang et al., 2024 note “the immense volume of training data necessitated for effective LM performance, surpassing—by orders of magnitude—the linguistic input received during human language acquisition” Zhuang, Fedorenko, and Andreas (2024). People, by contrast, are very good at wringing new information from familiar stories.110110110Right now, models display a striking dissimilarity with people in what happens when they are repeatedly exposed to the same data, either directly, or indirectly through iterated training on synthetic data. For example, Chang et al., 2024 found that “LLMs trained with duplicated training data exhibit faster forgetting” Chang et al. (2024), and Shumailov et al., 2024, found “indiscriminate use of model-generated content in training causes irreversible defects in the resulting models, in which tails of the original content distribution disappear” Shumailov et al. (2024). On the other hand, a hallmark of human storytelling is enjoying—and getting new content out of—experiencing a story (created by other humans, even potentially by oneself) again and again. People have been studying religious texts, as in Biblical exegesis, for thousands of years, yet new ideas emerge. An individual may return to the Bible (continuing the same example) upwards of hundreds of times, and yet experience new insight. How can LLMs wring new insight from old data? This stark difference in behaviour is worth further consideration.
@@
One strategy would be to bring various representations the LLM uses into greater alignment with human structures and processes. There is evidence that humans can do quite a lot with very little data (e.g., a broad spectrum of poverty-of-the-stimulus-related arguments and processes Zhuang, Fedorenko, and Andreas (2024), including fast-mapping Gelman and Brandone (2010), and in experimental scenarios testing the distributional hypothesis McDonald and Ramscar (2001)).111111111Likewise, LLMs can sometimes do a lot with a little, such as in lexinvariant LMs Huang et al. (2023b).
@@
Furthermore, good model vocabularies may already be converging on being somewhat human-like. Symmetrically, the tokens themselves point to commonalities between the entropy-minimizing objective functions often used in tokenization and the forces that shape human language. The hundreds of thousands of years Fedorenko, Piantadosi, and Gibson (2024)112112112E. Fedorenko is an author on both Fedorenko et al.,, 2024 Fedorenko, Piantadosi, and Gibson (2024) and Zhuang et al.,, 2024 Zhuang, Fedorenko, and Andreas (2024), which seem to make arguments that are at least partially contradictory. We assume in good faith there must be a perspective (presumably E.F.’s) from which the arguments are entirely consistent, and take both papers at face value for the purposes of this project. in which only people spoke language, during which pressures of efficiency seem to have shaped its development at multiple levels, surely evolved structures we can now take advantage of in our models.
@@
We argue that the de facto approaches in each of those areas are worth revisiting.113113113The use of special tokens (e.g., SEP) in LLMs is out of scope for this paper but intriguing for future work, as is the impact of embedding vector dimension and other implementation choices. More broadly, now that the efficacy of LLM architecture has been established, we should revisit our prior underconstrained implementation choices, such as the relevant objective functions that shape the model and the functions that conjoin its components.114114114e.g., operations involving Query, Key, and Value vectors in the attention blocks, or applications of softmax.
@@
We contend that LLMs validate and isolate the distributional hypothesis. Specifically, they show the power of syntagmatic information for producing minimally viable language performance (when coupled with sufficient architecture). We hypothesize that LLMs bootstrap limited meaning, potentially including abstractions and world knowledge, from distributional patterns: more specifically, that while LLM’s start with syntagmatics, they create further structure leading to limited semantic meaning, the ceiling of which is what can be conveyed by purely distributional linguistic information and diegetic approximations composed thereof.115115115Harris’ purely linguistic meaning Sahlgren (2008). The lack of hard boundary between world knowledge and linguistic knowledge makes this feat possible but it also engenders the problems that come from fuzzing up grammaticality and truth, or frequency and truth Zimmerman et al. (2024); Gong et al. (2020),116116116Gong et al., 2020, found that in some situations, word embeddings were “biased towards word frequency”, with the embeddings of high- and low-frequency words lying in different areas of the embedding space, and the embedding of a rare word and a popular word being physically distant despite their semantic similarity. Within the embedding space, rare and popular words behaved differently, with popular words having semantically relevant neighbors, and rare words being overwhelmingly surrounded by other, semantically-unrelated, rare words Gong et al. (2020). (They foresee these structures as causing downstream performance issues, so they propose FRequency AGnostic word Embedding (FRAGE) as a solution Gong et al. (2020).) which are amplified by the choice of next token prediction as objective function.117117117And by other, more specific, architectural choices, such as how attention is implemented. These choices are related, since the attention design is in the service of generating plausible output tokens, so in the context of this paper, we are referring to the broad goal of getting the model to generate plausible tokens, not only to specific portions of the model that might be most directly technically implicated, like the decoder (if there is one), the output layer, or the final probability distribution over the vocabulary. We propose that experiential grounding enhances the depth of meaning through othermodal reference for both people and LLMs118118118Recall Wolfe and Caliskan, 2022 Wolfe and Caliskan (2022) and Zhuang et al., 2024 Zhuang, Fedorenko, and Andreas (2024). and therefore represents a promising avenue for their continued improvement.
@@
Importantly, a model’s gnogeography119119119Recall that gnogeography is the landscape of knowledge; a being’s mind-and-body, their architecture, in conjunction with what information they have taken in. can be expanded through diegetic approximations Zimmerman et al. (2024); Brown et al. (2020); Chang et al. (2024); Mahowald et al. (2024). Meaning that, with careful structuring of training data and model architecture (e.g., batch size, number of training steps), LLMs do seem to be able to acquire additional world knowledge Chang et al. (2024), implying that there is space between the world knowledge necessary for MVP language and what it is possible to know through the medium of language. Given that current LLMs clearly possess MVP language but are missing much human-familiar knowledge and common sense, they are somewhere within that space Zimmerman et al. (2024); Mahowald et al. (2024). The entirety of that space is the upper bound of their gnogeography.
@@
This is crucial to remember in the face of AI hype, one core promise of which is “we have a machine that can talk, therefore we have a machine that can think”. In fact, LLMs demonstrate the plausibility of a machine that can solely talk; while a machine may be able to both talk and think, the ‘therefore’ between those abilities is not justified.120120120“Talk” and “think” are shorthand. Of course, talking involves thinking, and thinking can be scaffolded by talking. See previous sections, especially Sec. 2.3, where we referenced stochastic parrots; we do think MVP language as implemented in LLMs involves more understanding than would be granted by that framework.
@@
The size and shape of the upper bound LLM gnogeography is unknown.121121121The hypothetical ceiling being everything that can be encoded in language. While this is exciting, it also means it is plausible that their gnogeography is incompatible with what would be required for many of the tasks we want LLMs to perform. This warning also applies to context-specific modifications to the model such as alignment and fine-tuning: strategies that do not attempt to reconcile the innate gnogeographic constraints of the model with the intended task have a flimsy foundation on which to build success. As long as generating plausible text is the goal for the model, we should expect plausible text rather than accurate text to be the output. That those two can often coincide, while alluring and often convenient, is fundamentally incidental.122122122We think this framework for thinking about LLMs can be helpful to people trying to decide how to use and study them. If an LLM is deployed as part of e.g., a transcription pipeline, the potential for unwanted hallucination can and should be assumed by default Koenecke et al. (2024), unless part of the pipeline’s design is specifically intended to counteract not just the symptomatic output but the causes - and even then, caution is necessary.
5 Post-Script: what’s next
@@
The architectural choice of the token, while foundational, may be on its way out for some AI applications. By hardwiring a specific, language-centric unit of cognition, tokens impose a form of instantiated linguistic relativism – a commitment to the primacy of language for thought. As discussed above, the tokens themselves are determinative of distributional and organizational unit, and carry consequences such as ‘flattening’. However, the commitment to linguistic primacy is even more pervasive – it manifests not only in tokens’ role as semantic primitives and as the units governing the distributional information perceptible to the model, but also in their function in Chain-of-Thought-like (CoT) reasoning123123123What exactly ‘reasoning’ models are doing – whether via CoT or other means – remains an open question. We believe there are compelling reasons to think the gnogeographic ceiling of deep learning could be far higher than that attested by current prominent LLMs Zimmerman (2025). However, model behavior is notoriously difficult to interpret, so caution is warranted. For concision, here we assume CoT induces something beyond MVP linguistic performance – a substantive rather than stylistic difference. However, it is possible that ‘reasoning’ LLMs may more aptly be described as reweighting statistical patterns under varying conditions rather than truly reasoning about new tasks. and their direct relationship with computational resource allocation during inference (e.g., the correspondence between a single forward pass and a single token during generation) Gladstone et al. (2025); Wang et al. (2025). Historically, more open-ended architectures have often unlocked greater performance, suggesting that the degree of cognition currently outsourced to tokens as an architectural choice might be better left to the model itself to learn Zimmerman (2025).
@@
This plausible-text-generating architecture makes the emergence of ‘reasoning’ all the more puzzling. The primary objective for a model in Next-Token-Prediction (NTP) is to create plausible text based on prior text – effectively treating frequency as truth. The better-framed question, then, is not ‘why do models hallucinate?’ but rather, why does something that looks like reasoning ever emerge from this setup? Can the local task of predicting the next token genuinely be extended into something like thinking?
@@
We speculate that the answer lies in the expanding scope of knowledge required by LLM tasks and thus incorporated into LLM training objectives. NTP more-or-less suffices for human-like text generation. However, the shift towards ‘reasoning’ – epitomized by CoT prompting – demands a broader scope. The task extends beyond MVP linguistic competence to solving multipart logical problems that invoke narrative, social rules, specialized techniques, and extensive world knowledge. Consequently, effective CoT and instruction tuning must engage larger linguistic structures than NTP alone. We view CoT as a possible extension of the NTP objective, broadening the model’s perceptible range of knowledge and raising the ceiling of its gnogeography.124124124Essentially, as tasks demand more from the information encoded in language, models must become increasingly aware of linguistic structures not salient within NTP alone (at least generally, many of these structures would seem to be larger than that local scope). Thus, as ‘reasoning’ LLMs improve, they increasingly demonstrate the feasibility of language as a cognitive foundation, one strategy for how it is possible to think. Caveats remain: the boundary between world knowledge and linguistic knowledge is unclear, perhaps non-existent, and NTP may foster some non-local abstractions Dumas et al. (2025); Cui et al. (2024); Zimmerman (2025); Hewitt et al. (2024); Digutsch and Kosinski (2023); Huang et al. (2023b). Relationships among task design, implementation, model capacity, and data remain only loosely understood – but in broad strokes, this is the emerging picture.
@@
Simultaneously, in refining what they want LLMs to do, computer scientists are often rediscovering principles from linguistic pragmatics. For instance, the CoT principles of efficacy, efficiency, brevity, and sufficiency directly echo Grice’s maxims of quantity, quality, relation, and manner Cheng et al. (2025); Grice (1989). Similarly, a Gricean logic is evident in efforts to incentivize ‘truthfulness’ by teaching a model to abstain when uncertain Wei et al. (2025),125125125‘Truthfulness’ is a loaded concept mostly out of scope for this discussion; we note ‘truth’ is – at the very least – unlikely to be a property found within the four corners of the text, loosely paraphrasing Joseph Weizenbaum. and in instruction tuning. These maxims emerge at the scale of conversation and multi-utterance discourse – scales that ‘reasoning’ models are increasingly forced to navigate. These resonances demonstrate that computer scientists could look to linguistics (and psychology and cognitive science) for inspiration, and in particular, for strategies which could enhance model performance – and vice versa!126126126Regardless of their architectural suitability, tokens remain invaluable to linguists: they are configurable, language-related structures generated by the most performant computational models of language yet developed. We plan to explore their use in linguistics further in future work (and hope others will be doing the same)!